# Weyl law for 1-cycles

### v1

@@ 4.3. Parametric isoperimetric inequality in compact manifolds
Now we proceed to define $R_{Q}(x,t)$. Let $q$ be the center of $Q$. Consider the map $I:\tilde{B}\to\tilde{B}\cap\partial Q$ which maps each point $z\in\tilde{B}$ to the point $I(z)$ in the line through $q$ and $z$ and in $\partial\tilde{B}$ which is furthest from $q$ (and hence it is in $\tilde{B}\cap\partial Q$). Let $I_{t}(z)$ be a constant speed parametrization of the segment from $z$ to $I(z)$. This yields a Lipschitz homotopy $I:\tilde{B}\times[0,1]\to\tilde{B}$ with $I(z,0)=z$ and $I(z,1)=I(z)$ for each $z\in\tilde{B}$. This is analog to the map $H$ previously defined but replacing $Q$ by $\tilde{B}$ and $P$ by $q$. We proceed to replace $R_{Q}(x)$ by the collection of segments $\Xi(\partial R_{Q}(x),I(\partial R_{Q}(x)))$ in the following way
@@
Observe that the number of segments in $\Xi(I_{t}(F^{\prime}_{Q}(x)-F_{Q}(x)),F^{\prime}_{Q}(x)-F_{Q}(x))$ is equal to $\operatorname{\mathbf{M}}((F^{\prime}_{Q}(x)-F_{Q}(x))\llcorner\operatorname{int}(\tilde{B}))$ because points in $\partial\tilde{B}\cap\partial Q$ are fixed by $I$ and $(F^{\prime}_{Q}(x)-F_{Q}(x))\llcorner(\partial\tilde{B}\setminus\partial Q)=0$. As $\operatorname{\mathbf{M}}(F_{Q}(x)\llcorner\operatorname{int}(Q))\leq\operatorname{\mathbf{M}}_{Q}$ and $\operatorname{\mathbf{M}}(F^{\prime}_{Q}(x)\llcorner\operatorname{int}{Q})\leq\operatorname{\mathbf{M}}_{Q}+p+1$, we deduce that
@@
From our definition also follows that
$\operatorname{\mathbf{M}}(\partial\Xi(I_{1}(F^{\prime}_{Q}(x)-F_{Q}(x)),F^{\prime}_{Q}(x)-F_{Q}(x))\llcorner\partial Q)$ $\leq\operatorname{\mathbf{M}}((F_{Q}-F_{Q}^{\prime}(x))\llcorner\operatorname{int}(\tilde{B}))$
@@
Theorem 4.19.
Let $(M,g)$ be an $n$-dimensional compact triangulated Riemannian manifold with corners. Let $P\subseteq M$ be an $m$-dimensional subcomplex of $M$. There exists a constant $C(P)$ depending only on $P$ such that the following holds. Let $F:X^{p}\to\mathcal{Z}_{0}(P)$ be a contractible family of 00-cycles on $P$ and denote
Then there exists a continuous family $G:X^{p}\to\mathcal{I}_{1}(M)$ such that
  1. (1)
    $\partial G(x)=F(x)$.
  2. (2)
Proof.
By induction on $m=\dim(P)$. If $m=1$, let $C(P)$ be the length of the $1$-dimensional polyhedron $P$. Using [12][Proposition 3.5], we can construct a family $G:X\to\mathcal{I}_{1}(P)$ such that $\partial G(x)=F(x)$ (more details on how to do this are provided in the proof of the inductive step) and observe that as we are taking coefficients in $\mathbb{Z}_{2}$ it holds $\operatorname{\mathbf{M}}(G(x))\leq C(P)$, verifying the required estimate. Suppose the result is true for $m-1$ and let $P$ be an $m$-dimensional subcomplex of $M$. By [12][Proposition 3.5], for every $\eta>0$ there exists a continuous family $\tilde{G}:X\to\mathcal{I}_{1}(M)$ such that
@@
Take $\eta<<p^{\frac{m-1}{m}}$ and use Theorem 3.16 to find a $\delta$-localized family $G^{\prime}:X^{p}\to\mathcal{I}_{1}(M)$ such that $\mathcal{F}(G^{\prime}(x),\tilde{G}(x))\leq\eta$ and
$\sup_{x\in X}\{\operatorname{\mathbf{M}}(F^{\prime}(x))\}\leq\operatorname{\mathbf{M}}_{0}$
where $F^{\prime}=\partial G^{\prime}$ (here we are using that $F^{\prime}$ is a family of 00-cycles). Notice that $\delta$ can also be taken to be arbitrarily small. As
@@
by Theorem 3.48 we can construct a family $G^{\prime\prime}:X^{p}\to\mathcal{I}_{1}(M)$ such that $\partial G^{\prime\prime}(x)=F(x)-F^{\prime}(x)$ and $\operatorname{\mathbf{M}}(G^{\prime\prime}(x))\leq C(p)\eta$. Therefore if $G_{1}(x)=G^{\prime}(x)+G^{\prime\prime}(x)$ then $\partial G_{1}(x)=F(x)$, $G^{\prime\prime}(x)$ has very small mass, $G^{\prime}$ and $F^{\prime}=\partial G^{\prime}$ are $\delta$-localized and $F^{\prime}$ has mass bounded by $\operatorname{\mathbf{M}}_{0}$. Therefore, it is enough to prove the theorem for the family $F^{\prime}$ by deforming its $\delta$-localized filling $G^{\prime}$ cell by cell of $P$.
@@
and
@@
Let $\hat{G}(x)=\overline{G}(x)-\sum_{j=1}^{J}\overline{G}_{j}(x)$ and $\hat{F}=\partial\hat{G}$. Let $P^{\prime}$ denote the $(m-1)$-skeleton of $P$. Observe that as $\operatorname{spt}(\hat{G}(x))\subseteq P^{\prime}$, $\hat{F}$ is a contractible family of 00-cycles supported in $P^{\prime}$ verifying
@@ 5. Weyl Law for the Volume Spectrum
In this section we prove the Weyl Law for the Volume Spectrum for $1$-cycles (Theorem 1.1), using the Parametric Coarea Inequality and the Parametric Isoperimetric Inequality stated in the introduction. Fix an $n$-dimensional compact Riemannian manifold $(M^{n},g)$, possibly with boundary. We know that the Weyl law holds for contractible domains of $\mathbb{R}^{n}$. The previous means that there exists a universal constant $\alpha(n)$ such that if $U$ is a compact contractible domain in $\mathbb{R}^{n}$ with Lipschitz boundary (or Lipschitz domain according to the terminology in [15]) then
$\lim_{p\to\infty}\omega_{p}^{1}(U)p^{-\frac{n-1}{n}}=\alpha(n)\operatorname{Vol}(U)^{\frac{1}{n}}.$
In [15] it was also showed that for a compact $n$-manifold $(M,g)$,
@@
$\limsup_{p\to\infty}\omega_{p}^{1}(M,g)p^{-\frac{n-1}{n}}\leq\alpha(n,1)\operatorname{Vol}(M,g)^{\frac{1}{n}}.$
With that purpose, we want to construct for each $p\in\mathbb{N}$ an “almost optimal” (at least asymptotically up to an error of the order $o(p^{\frac{n-1}{n}})$) $p$-sweepout of $(M,g)$ to obtain good upper bounds for $\omega_{p}^{1}(M,g)$.
Fix $\varepsilon>0$. Using [3][Theorem 1.2 and 1.3], consider a $(1+\frac{\varepsilon}{2})$-triangulation of $M$ via simplices $\{Q_{i}\}_{1\leq i\leq N}$ (see Definition 2.2). This implies that each $Q_{i}$ is $(1+\frac{\varepsilon}{2})$-bilipschitz diffeomorphic to a compact contractible domain $U_{i}$ in $\mathbb{R}^{n}$ (more precisely, each $U_{i}$ is a linear $n$-simplex). Denote $\Phi_{i}:Q_{i}\to U_{i}$ such bilipschitz diffeomorphisms. We can translate the $U_{i}$’s if necessary and connect each $U_{i}$, $i>2$ with $U_{1}$ via a tiny tube $T_{i}$ obtaining a compact contractible domain $U=U_{1}\cup\bigcup_{i=2}^{N}U_{i}\cup T_{i}$ with Lipschitz boundary such that $\operatorname{Vol}(U)<(1+\varepsilon)^{n}\operatorname{Vol}(M,g)$.
Now fix $p\in\mathbb{N}$. By definition of $p$-width, we can find a $p(n-1)$-dimensional cubical complex $X(p)$ and a $p$-sweepout $\tilde{F}^{p}:X(p)\to\mathcal{I}_{1}(U,\partial U)$ of $U$ such that
$\sup_{x\in X(p)}\operatorname{\mathbf{M}}(\tilde{F}^{p}(x))\leq(1+\varepsilon)\omega^{1}_{p}(U).$
@@
(12) $\operatorname{\mathbf{M}}(F^{p}(x))\leq(1+\varepsilon)\operatorname{\mathbf{M}}(\tilde{F}^{p}(x))\leq(1+\varepsilon)^{2}\omega_{p}^{1}(U).$
However, $F^{p}$ is not a family of relative cycles in $(M,\partial M)$ because it may have arbitrarily large boundary in $\bigcup_{i=1}^{N}\partial Q_{i}$ which is not contained in $\partial M$. Denote $V_{m}=\bigcup_{i=1}^{m}Q_{i}$. To fix the previous, we are going to construct inductively a short $p$-sweepout $G_{m}$ of $(V_{m},\partial V_{m})$ which will finally provide a short $p$-sweepout $G=G_{N}$ of $(M,\partial M)$. To be precise, we want $G_{m}$ to obey the following mass bounds
@@
In order to define the $G_{m}$, we will need to apply Theorem 1.4 to each $F^{p}_{i}$. For that purpose we fix $0<\alpha<\frac{1}{n-1}-\frac{1}{n}$. We obtain a new $p$-sweepout $\overline{F}^{p}_{i}:X(p)\to\mathcal{Z}_{1}(Q_{i},\partial Q_{i})$ which is induced by a family of absolute chains $\overline{F}^{p}_{i}:X(p)\to\mathcal{I}_{1}(Q_{i})$ with $\operatorname{spt}(\partial\overline{F}^{p}_{i}(x))\subseteq\partial Q_{i}$ such that
@@
Denote $A_{m}=\partial V_{m}\cap\partial Q_{m+1}$. We want to glue the families $G_{m}^{p}$ and $\overline{F}_{m+1}^{p}$ along $A_{m}$ to get a $p$-sweepout of $V_{m}\cup Q_{m+1}=A_{m+1}$. We know that $\partial G_{m}^{p}(x)\llcorner A_{m}$ and $\partial\tilde{F}^{p}_{m+1}(x)\llcorner A_{m}$ are $p$-sweepouts of $A_{m}$ by 00-cycles. Despite they may not coincide (preventing us from gluing $G_{m}^{p}$ and $F_{m+1}^{p}$ directly), their difference $\eta^{p}_{m}:X(p)\to\mathcal{Z}_{0}(A_{m},\partial A_{m})$ is a contractible family of 00-cycles in $(A_{m},\partial A_{m})$ with
@@
$\leq(1+\varepsilon)^{3}\alpha(n)\operatorname{Vol}(M,g)^{\frac{1}{n}}$
by (12). As that holds for every $\varepsilon>0$ we deduce

### v2

@@ 4.3. Parametric isoperimetric inequality in compact manifolds
Now we proceed to define $R_{Q}(x,t)$. Let $q$ be the center of $Q$. Consider the map $I:\tilde{B}\to\tilde{B}\cap\partial Q$ which maps each point $z\in\tilde{B}$ to the point $I(z)$ in the line through $q$ and $z$ and in $\partial\tilde{B}$ which is furthest from $q$ (and hence it is in $\tilde{B}\cap\partial Q$). Let $I_{t}(z)$ be a constant speed parametrization of the segment from $z$ to $I(z)$. This yields a Lipschitz homotopy $I:\tilde{B}\times[0,1]\to\tilde{B}$ with $I(z,0)=z$ and $I(z,1)=I(z)$ for each $z\in\tilde{B}$. This is analog to the map $H$ previously defined but replacing $Q$ by $\tilde{B}$ and $P_{0}$ by $q$. We proceed to replace $R_{Q}(x)$ by the collection of segments $\Xi(\partial R_{Q}(x),I(\partial R_{Q}(x)))$ in the following way
@@
Observe that the number of segments in $\Xi(I_{t}(F^{\prime}_{Q}(x)-F_{Q}(x)),F^{\prime}_{Q}(x)-F_{Q}(x))$ is equal to $\operatorname{\mathbf{M}}((F^{\prime}_{Q}(x)-F_{Q}(x))\llcorner\operatorname{int}(Q))$ because points in $\partial\tilde{B}\cap\partial Q$ are fixed by $I$ and $(F^{\prime}_{Q}(x)-F_{Q}(x))\llcorner(\partial\tilde{B}\setminus\partial Q)=0$. As $\operatorname{\mathbf{M}}(F_{Q}(x)\llcorner\operatorname{int}(Q))\leq\operatorname{\mathbf{M}}_{Q}$ and $\operatorname{\mathbf{M}}(F^{\prime}_{Q}(x)\llcorner\operatorname{int}{Q})\leq\operatorname{\mathbf{M}}_{Q}+p+1$, we deduce that
@@
From our definition, it also follows that
$\operatorname{\mathbf{M}}(\partial\Xi(I_{1}(F^{\prime}_{Q}(x)-F_{Q}(x)),F^{\prime}_{Q}(x)-F_{Q}(x))\llcorner\partial Q)$ $\leq\operatorname{\mathbf{M}}((F_{Q}(x)-F_{Q}^{\prime}(x))\llcorner\operatorname{int}(Q))$
@@
Proof of Theorem 4.18.
By induction on $m=\dim(P)$. If $m=1$, let $C(P)$ be the length of the $1$-dimensional polyhedron $P$. Using [GL22][Proposition 3.5], we can construct a family $G:X\to\mathcal{I}_{1}(P)$ such that $\partial G(x)=F(x)$ (more details on how to do this are provided in the proof of the inductive step) and observe that as we are taking coefficients in $\mathbb{Z}_{2}$ it holds $\operatorname{\mathbf{M}}(G(x))\leq C(P)$, verifying the required estimate. Suppose the result is true for m1m-1 and let $P$ be an $m$-dimensional Riemannian polyhedron. By [GL22][Proposition 3.5], for every $\eta>0$ there exists a continuous family $\tilde{G}:X\to\mathcal{I}_{1}(M)$ such that
@@
Take $\eta<<p^{\frac{m-1}{m}}$ and use Theorem 3.13 to approximate $\tilde{G}$ by a $\delta$-localized family $G^{\prime}:X^{p}\to\mathcal{I}_{1}(M)$ such that $\mathcal{F}(G^{\prime}(x),\tilde{G}(x))\leq\eta$ and if $F^{\prime}=\partial G^{\prime}$ then
$\sup_{x\in X}\{\operatorname{\mathbf{M}}(F^{\prime}(x))\}\leq\operatorname{\mathbf{M}}_{0}.$
In the previous, we used that $F^{\prime}$ is a family of 0-cycles. Notice that $\delta$ can also be taken to be arbitrarily small. As
@@
by Theorem 3.45 we can construct a family $G^{\prime\prime}:X^{p}\to\mathcal{I}_{1}(M)$ such that $\partial G^{\prime\prime}(x)=F(x)-F^{\prime}(x)$ and $\operatorname{\mathbf{M}}(G^{\prime\prime}(x))\leq\varepsilon$ for a small $\varepsilon>0$ provided $\eta$ is small enough. Therefore if $G_{1}(x)=G^{\prime}(x)+G^{\prime\prime}(x)$ then $\partial G_{1}(x)=F(x)$, $G^{\prime\prime}(x)$ has very small mass, $G^{\prime}$ and $F^{\prime}=\partial G^{\prime}$ are $\delta$-localized and $F^{\prime}$ has mass bounded by $\operatorname{\mathbf{M}}_{0}$. Therefore, it is enough to prove the theorem for the family $F^{\prime}$ by deforming its $\delta$-localized filling $G^{\prime}$ cell by cell of $P$.
@@
(plus a small error in terms of $\delta$ which will converge to 0 as we choose $\delta\to 0$) and
@@
Let $\hat{G}(x)=\overline{G}(x)-\sum_{j=1}^{J}\overline{G}_{j}(x)$ and $\hat{F}=\partial\hat{G}$. Let $P^{\prime}$ denote the (m1)(m-1)-skeleton of $P$. Observe that as $\operatorname{spt}(\hat{G}(x))\subseteq P^{\prime}$, $\hat{F}$ is a contractible family of 0-cycles supported in $P^{\prime}$ verifying
@@ 5. Weyl Law for the Volume Spectrum
In this section we prove the Weyl Law for the Volume Spectrum for $1$-cycles (Theorem 1.1), using the Parametric Coarea Inequality and the Parametric Isoperimetric Inequality (Theorem 1.4). Fix an $n$-dimensional compact Riemannian manifold $(M^{n},g)$, possibly with boundary (in fact, this proof works for almost $1$-Lipschitz triangulable piecewise smooth Riemannian manifolds, see Definition 2.7). We know that the Weyl law holds for contractible domains of $\mathbb{R}^{n}$. The previous means that there exists a universal constant $\alpha(n)$ such that if $U$ is a compact contractible domain in $\mathbb{R}^{n}$ with Lipschitz boundary (or Lipschitz domain according to the terminology in [LMN]) then
$\lim_{p\to\infty}\omega_{p}^{1}(U,\partial U)p^{-\frac{n-1}{n}}=\alpha(n)\operatorname{Vol}(U)^{\frac{1}{n}}.$
In [LMN] it was also showed that for a compact $n$-manifold $(M,g)$,
@@
$\limsup_{p\to\infty}\omega_{p}^{1}(M,g)p^{-\frac{n-1}{n}}\leq\alpha(n)\operatorname{Vol}(M,g)^{\frac{1}{n}}.$
With that purpose, we want to construct for each $p\in\mathbb{N}$ an “almost optimal” (at least asymptotically up to an error of the order $o(p^{\frac{n-1}{n}})$) $p$-sweepout of $(M,g)$, to obtain good upper bounds for $\omega_{p}^{1}(M,g)$.
Fix $\varepsilon>0$. Using [Bowditch][Theorem 1.2 and 1.3], consider a $(1+\frac{\varepsilon}{2})$-bilipschitz smooth triangulation of $M$ via simplices $\{Q_{i}\}_{1\leq i\leq N}$ (see Definition 2.4). This implies that each $Q_{i}$ is $(1+\frac{\varepsilon}{2})$-bilipschitz diffeomorphic to a compact contractible domain $U_{i}$ in $\mathbb{R}^{n}$ (more precisely, each $U_{i}$ is a linear $n$-simplex). Denote $\Phi_{i}:Q_{i}\to U_{i}$ such bilipschitz diffeomorphisms. We can translate the $U_{i}$’s if necessary and connect each $U_{i}$, $i>1$ with $U_{1}$ via a tiny tube $T_{i}$ obtaining a compact contractible domain $U=U_{1}\cup\bigcup_{i=2}^{N}U_{i}\cup T_{i}$ with Lipschitz boundary such that $\operatorname{Vol}(U)<(1+\varepsilon)^{n}\operatorname{Vol}(M,g)$.
Now fix $p\in\mathbb{N}$. By definition of $p$-width, we can find a p(n1)p(n-1)-dimensional cubical complex $X(p)$ and a $p$-sweepout $\tilde{F}^{p}:X(p)\to\mathcal{I}_{1}(U,\partial U)$ of $U$ such that
$\sup_{x\in X(p)}\operatorname{\mathbf{M}}(\tilde{F}^{p}(x))\leq(1+\varepsilon)\omega^{1}_{p}(U,\partial U).$
@@
(12) $\operatorname{\mathbf{M}}(F^{p}(x))\leq(1+\varepsilon)\operatorname{\mathbf{M}}(\tilde{F}^{p}(x))\leq(1+\varepsilon)^{2}\omega_{p}^{1}(U,\partial U).$
However, $F^{p}$ is not a family of relative cycles in $(M,\partial M)$ because it may have an arbitrarily large boundary supported in $\bigcup_{i=1}^{N}\partial Q_{i}$ which is not contained in $\partial M$. Denote $V_{m}=\bigcup_{i=1}^{m}Q_{i}$. To fix the previous, we are going to construct inductively a short $p$-sweepout $G_{m}$ of $(V_{m},\partial V_{m})$ which will finally provide a short $p$-sweepout $G=G_{N}$ of $(M,\partial M)$. To be precise, we want $G_{m}$ to obey the following mass bounds
@@
In order to define the $G_{m}$, we will need to apply Theorem 1.4 to each $F^{p}_{i}$ (this can be done because the theorem holds for simplices as they are PL submanifolds of $\mathbb{R}^{n}$, see [StaCoarea]). For that purpose we fix $0<\alpha<\frac{1}{n-1}-\frac{1}{n}$. We obtain a new $p$-sweepout $\overline{F}^{p}_{i}:X(p)\to\mathcal{Z}_{1}(Q_{i},\partial Q_{i})$ which is induced by a family of absolute chains $\overline{F}^{p}_{i}:X(p)\to\mathcal{I}_{1}(Q_{i})$ with $\operatorname{spt}(\partial\overline{F}^{p}_{i}(x))\subseteq\partial Q_{i}$ such that
@@
Denote $A_{m}=\partial V_{m}\cap\partial Q_{m+1}$. We want to glue the families $G_{m}^{p}$ and $\overline{F}_{m+1}^{p}$ along $A_{m}$ to get a $p$-sweepout of $V_{m}\cup Q_{m+1}=A_{m+1}$. We know that $\partial G_{m}^{p}(x)\llcorner A_{m}$ and $\partial\tilde{F}^{p}_{m+1}(x)\llcorner A_{m}$ are $p$-sweepouts of $A_{m}$ by 0-cycles. Despite they may not coincide (preventing us from gluing $G_{m}^{p}$ and $F_{m+1}^{p}$ directly), their difference $\eta^{p}_{m}:X(p)\to\mathcal{Z}_{0}(A_{m},\partial A_{m})$ is a contractible family of 0-cycles in $(A_{m},\partial A_{m})$ with
@@
$\leq(1+\varepsilon)^{n+2}\alpha(n)\operatorname{Vol}(M,g)^{\frac{1}{n}}$
by (12). As that holds for every $\varepsilon>0$, we deduce
@@
Department of Mathematics, Rice University, Houston, TX 77005, USA
Email address: bruno.staffa@rice.edu